OpenStudy (superhelp101):

The grimsley inn has 250 rooms and the standard rate is $150 per night per room. During the non-holiday season, 100 rooms are booked each day on average. But, for each $20 price reduction, 25 more rooms will be rented. What non-holiday price should be advertised?

OpenStudy (superhelp101):

@aum @freckles

OpenStudy (superhelp101):

@jim_thompson5910

OpenStudy (aum):

Assume price reduction of n * 20. Rate per night = 150 - 20n Number of rooms booked = 100 + 25n Total revenue R = (150 - 20n) * (100 + 25n) Maximize R. Find dR/dn, equate to zero and solve for n.

ganeshie8 (ganeshie8):

or simply find the vertex of quadratic

ganeshie8 (ganeshie8):

they tricky part is setting up revenue function as u can see

OpenStudy (superhelp101):

wait, so what do we do with the 250 rooms

OpenStudy (aum):

That is a constraint on the max rooms that can be rented.

OpenStudy (aum):

100 + 25n <= 250

OpenStudy (superhelp101):

lol what does that mean?

OpenStudy (aum):

n <= 6.

OpenStudy (superhelp101):

is it possible to have 250 instead of 25 ?

OpenStudy (aum):

When finding the "n" that gives maximum revenue, if n exceeds 6, then you have to stop at n = 6.

OpenStudy (superhelp101):

okay it's making a bit more sense. But I am not completing understanding how we got 100+25n

OpenStudy (aum):

"for each $20 price reduction, 25 more rooms will be rented." If price reduces by 20*n, then 25*n more rooms will be rented. "More" here refers to more than average number of rooms rented when the rate is standard. That is 100 rooms + 25*n rooms.

OpenStudy (superhelp101):

oh okay

OpenStudy (superhelp101):

Thank you!

OpenStudy (aum):

You are welcome.